An ideal gas undergoes a mechanically reversible

process (constant pressure, constant

temperature and adiabatic process. The gas entering a T1=650K and P1=10bar decreases its temperature at constant pressure where V2=2.91x10^-3 m^-3. Then, it went to isothermal process to decrease its pressure. Finally, the gas returns to its initial state. Take CP = 7/2 and CV = 5/2R.


a. T2, P3, V1and V3

b.Q,W,ΔU and ΔH



1
Expert's answer
2022-02-22T08:18:01-0500

T1=650K


P1= 10bar


V2=2.91x10-3


Constant P and T


V1=?


Pdv= nRTPdv=nRT


10(V2-V1) = 8.314×650


2.91x10-2 -10v1 =5404.1


V1=5.4×103